What is the smallest sample size needed so that a sampling distribution of a sample proportion can be claimed approximately normal from a population with a proportion of p=15%?

(Round to the next whole number)
1
Expert's answer
2020-08-03T18:44:07-0400

The proportion of a random sample has a sampling distribution whose shape can be approximated by a normal model if "np\geq10" and "np(1-p)\geq10."

Given "p=0.15"


"0.15n\geq10""0.15(1-0.15)n\geq10"

"n\geq\dfrac{10}{0.15}""n\geq\dfrac{10}{0.15\cdot0.85}"

Since "n" is integer


"n\geq79"

The smallest sample size is "79."